Consider the following two functions;

1. f: R-R defined by f(x) = 4x-15.

2. g: RR-defined by f(x) = 15x3.

Prove that both f and g are one-to-one correspondence.

1
Expert's answer
2021-05-07T09:04:01-0400
  1. Given "f:R\rightarrow R" defined by "f(x)=4x-5"

Require to prove that "f" is a one-to-one correspondence.

To prove that "f" is a one-to-one correspondence, we need to prove that "f" is one-one and onto.


One - one:

Let "a,b\in R" such that "f(a)=f(b)"

We prove that "a=b"

Now "f(a)=f(b)\Rightarrow 4a-5=4b-5"

"\Rightarrow 4a=4b\Rightarrow a=b"

So, we proved that "f(a)=f(b)\Rightarrow a=b" for all "a,b\in R"

Therefore, "f" is one - one.


Onto:

Let "b\in R" (codomain)

We need to find "b\in R" (domain) such that "f(a)=b"

Now "f(a)=b\Rightarrow 4a-5=b\Rightarrow 4a=b+5\Rightarrow a=\frac{b+5}{4}\in R"

So, we proved that for each "b\in R" (codomain) "\exists a=\frac{b+5}{4}\in R" such that "f(a)=b"

Therefore, "f" is onto.

"f" is one - one and onto implies "f" is one - to - one correspondence.

Hence, "f" is a one-to-one correspondence.

  1. Given: "f:R\rightarrow R" defined by "f(x)=15x^3"

Require to prove that "f" is a one - to - one correspondence.


One - one:


Let "a,b\in R" such that "f(a)=f(b)"

Now "f(a)=f(b)\Rightarrow 15a^3=15b^3\Rightarrow a^3=b^3\Rightarrow a=b"

We proved that "f(a)=f(b)\Rightarrow a=b" for all "a,b\in R"

Therefore, "f" is one - one.


Onto:


Let "b\in R" (codomain)

We require to find "a\in R" (domain) such that "f(a)=b"

Now "f(a)=b\Rightarrow 15a^3=b\Rightarrow a^3=\frac{b}{15}\Rightarrow a=(\frac{b}{15})^{\frac{1}{3}}\in R"

So, we proved that for each "b\in R" "\exists a=(\frac{b}{15})^{\frac{1}{3}}\in R" such that "f(a)=b"

Therefore, "f" is onto.


Hence, "f" is a one - to one correspondence.